You plan on going on a 3 month vacation 8 months from now. You can pay $4,500 per month during the vacation, or you can pay $13,000 today. Which alternative should you take and how much does it save you in present value term if your investments earn 3.00% APR (compounded monthly)?
Answer #1

Value of $4500 for 3 months at the end of 8 months
n=3
pmt=4500
r=0.03/12
Compute PV using excel:
=PV(0.03/12,3,4500,0)
=$13,432.78

Discounting it 8 months back to present=$13,432.78/(1.0025^8)=$13,167.12

Hence one should pay $13,000 today as it is cheaper